OpenStudy (envy_seven):

I have an Algebra exam tomorrow. Can anyone just help me revise some of the basics?

OpenStudy (clalgee):

Algebra I I suspect? :)

OpenStudy (envy_seven):

Yes

OpenStudy (clalgee):

Alright then! Let's do this :) Give me 30 seconds.

OpenStudy (envy_seven):

Thanks! :)

OpenStudy (clalgee):

No problem. So I'm going to mix it up. I will ask you for definitions, formulas, and I'll need you to simplify a few equations okay?

OpenStudy (envy_seven):

Aight

OpenStudy (clalgee):

Alright, Let's start with Order of Operations. Explain the steps of PEMDAS. Then solve the equation using Order of Operations below.

OpenStudy (clalgee):

-2[3 -(-2)(6)]= ?

OpenStudy (envy_seven):

P = Parentheses E = Exponents? M = Multiplication D = Division A = Addition S = Subtraction

OpenStudy (envy_seven):

34?

OpenStudy (envy_seven):

Wait actually...

OpenStudy (clalgee):

Great! You've mastered the understanding of PEMDAS But as for solving the equations.... You took a different route there

OpenStudy (clalgee):

1.) Parenthesis in PEMDAS Multiply (-2)(6)= ?

OpenStudy (envy_seven):

-12

OpenStudy (envy_seven):

-12 - 3 = 15 because of double negative?

OpenStudy (envy_seven):

3 - (-12) is how it should be written actually, right?

OpenStudy (clalgee):

2.) There are no exponents so we move multiplication. Multiply -12 by _____ ? What do you think?

OpenStudy (clalgee):

You got it :)

OpenStudy (envy_seven):

then -2 x 15 is -30?

OpenStudy (clalgee):

You got it again!

OpenStudy (clalgee):

Where did you get 34 from though?

OpenStudy (envy_seven):

-2[3 -(-2)(6)]= ? I did 3 - (-2) first for some reason

OpenStudy (clalgee):

Oh alright! That's fine, but just as long you knew what mistake you made :) So now, I'm going to ask you for definitions.

OpenStudy (clalgee):

What is absolute value? Give me an example.

OpenStudy (envy_seven):

The number between 0 and (for example, 5)?

OpenStudy (envy_seven):

And if it was -7, the absolute value would be 7?

OpenStudy (envy_seven):

So an example would be: 10 is the absolute value of 10 and -10